OpenStudy (anonymous):

how do you find the argument of z when expressing a complex number in trigonometriic functions

OpenStudy (anonymous):

i know that z=-2

OpenStudy (anonymous):

and i know z=r(cos theta+i sin theta)

OpenStudy (anonymous):

so it would be \[2(\cos \theta +i \sin \theta)\]

OpenStudy (anonymous):

but i need to know how to find theta which i know it says \[\tan^{-1} b/a\]

OpenStudy (anonymous):

but i dont know how to get b or a

OpenStudy (anonymous):

For purely imaginary or real numbers, you can easily pinpoint the argument using a diagram like the one below: